OpenStudy (kitkat1):

Part A: The sun produces 3.9 x 10^33 ergs of radiant energy per second. How many ergs of radiant energy does the sun produce in 2.45 x 10^5 seconds? Part B: Which is the more reasonable measurement of the distance between the tracks on a DVD: 7.4 x 10^-4 mm or 7.4 x 10^4 mm? Justify your answer.

OpenStudy (kitkat1):

Whats the answer?

OpenStudy (cwrw238):

A = 3.9 x 10^33 times 2.45 x 10^5 ergs

OpenStudy (cwrw238):

first work out 3.8 * 2.45 then 10^33 * 10 ^5 = 10 ^(33+5) = 10^38

OpenStudy (cwrw238):

then combine them

OpenStudy (cwrw238):

use ur calculator to do the first part

OpenStudy (kitkat1):

Okay how do I combine them?

OpenStudy (cwrw238):

well its just first part x 10^38

OpenStudy (kitkat1):

10X38 u mean?

OpenStudy (kitkat1):

3.8 X 2.45=9.31 what do i do next?

OpenStudy (cwrw238):

no first work out the 3.8 * 2.45 - that is the 'first part'

OpenStudy (cwrw238):

then its 9.31 x 10^38

OpenStudy (cwrw238):

thats the answer 9.31 x 10^38 ergs

OpenStudy (kitkat1):

I have allitle trouble with the 9.31 X 10^38 what does that equal?

OpenStudy (cwrw238):

I think thats what they want as the answer - its called scientific form. its huge! it has 39 digits in it:- 931 + 36 zeroes

OpenStudy (kitkat1):

But what about the rest?

OpenStudy (kitkat1):

no Actually 3.9 X 2.45=9.555

OpenStudy (kitkat1):

so the answer is 9.55 X 10^38?

OpenStudy (cwrw238):

oh ok - my mistake 9.555 x 10^38 is correct answer

OpenStudy (kitkat1):

Thank u very much what about the part B?

OpenStudy (kitkat1):

I need an answer for the part B

OpenStudy (cwrw238):

gap between traks on DVD is very small In scientific notation this has a negative power - so for example 2x10^-4 = 0.0002

OpenStudy (cwrw238):

so which one of the 2 options in B do you think is right one?

OpenStudy (kitkat1):

7.4 X 10^-4

OpenStudy (cwrw238):

yup

OpenStudy (kitkat1):

Thank u very much can u help me with just one more?

OpenStudy (cwrw238):

sorry gotta go right now

OpenStudy (kitkat1):

okay thank u for the help